OpenStudy (anonymous):

Solve: (u + 3)(u - 3) = 8u

OpenStudy (anonymous):

expand the brackets first and then bring the 8u from the right hand side to the left and then factor.

OpenStudy (anonymous):

and solve.

OpenStudy (anonymous):

what do you mean by expanding the brackets?

OpenStudy (anonymous):

do you know the foil method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use that.

OpenStudy (anonymous):

ok. Then i get u^2-9=8u

OpenStudy (anonymous):

yep. now bring the 8u to the left hand side.

OpenStudy (anonymous):

So now I have u^2-8u-9=0

OpenStudy (anonymous):

ok, now factorize again and solve.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so after i factorize, i get (u-1)(u+9)=0

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

now solve to get u.

OpenStudy (anonymous):

then the values of u would be -1 and 9, right?

OpenStudy (anonymous):

nope. u-1=0 u+9=0 now solve.

OpenStudy (anonymous):

my fault, 1 and -9. got the signs mixed up

OpenStudy (anonymous):

yes, that's correct.

OpenStudy (anonymous):

ok. thanks!

OpenStudy (anonymous):

no problem, medal thanks! lol